Refraction 1. Aray of light traveling through air is incident upon a sheet of crown glass 19.20 52) at an angle of 300 to the normal. What angle of refraction? Ans: is the 2. A ray of light passes into water from air at an angle of 50 to the lakes surface. What is the angle of refraction? What is the speed of light in the water Ans 28.9; 2.25 10sm/s 3. Calculate the s of light in a diamond. Ans: 1.24 x 108m/s 4. A block of unkown material is submerged in water. Light in the water is incident on the surface at an angle of 310. The angle ofrefraction is 27o. What is the speed of light in the unknown material? Ans: 1.20 x 108m/s 5. What is the critical angle for light ra passing from diamond (n 2.42) into air? Ans: 24.42 6. What is the critical angle for light ra passing from quartz (n 4) into water 1.33)? Ans: 59. 7. The critical angle for light rays passing from water into air is 48.750, what is the index of refraction of water? Ans: 1.33 is the thinnest film of a 1.40 refractive index coating on glass (n 1.50) for which destructive interference of the green component (500 n of an incident white light beam in air can take place by reflection? Ans: 178 nm Please show work. Thank you
Answer #1

We know from Snell's law

\frac{sin\theta_{1}}{sin\theta_{2}} = \frac{n_{2}}{n_{1}}

Therefore

\theta_{2} = sin^{-1}\bigg(\frac{n_1}{n_2} sin\theta_1\bigg) = 19.20^0